The cost of producing 4 units is $204.80. The cost of producing 8 units is $209.60. How much does it cost to produce 12 units?
Answer by checkley77(12844)

4x+y=204.80
-(8x+y=209.60) subtract
---------------------
-4x=-4.8
x=-4.8/-4
x=1.2 multiplier for each unit produced.
4*1.2+y=204.80
4.8+y=204.80
y=204.80-4.8
y=200 operating/overhead cost.
Proof:
8*1.2+200=209.60
9.6+200=209.60
209.60=209.60
12*1.2+200=14.4+200=$214.10 cost to produce 12 units.